OpenStudy (anonymous):

In LMN, is a right angle, LM = 76, and MN = 40. What is m

OpenStudy (anonymous):

Do you have a diagram?

OpenStudy (anonymous):

theres no diagram on my question test

OpenStudy (anonymous):

;(

OpenStudy (anonymous):

so the answer would be 76 times 40

OpenStudy (anonymous):

Wait, let me redo. 76 should be hypotenuse.

OpenStudy (anonymous):

28° 32° 40° 58°

OpenStudy (anonymous):

these are my possible answers

OpenStudy (anonymous):



OpenStudy (anonymous):

sin90/76 = sinL/40 sin90*40 = 40 40/76 = .5263157 sin-1(.5263157) = 31.7

OpenStudy (anonymous):

it was wrong

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Was it 58?

OpenStudy (anonymous):

it was 58 for people who are wondering